OpenStudy (anonymous):

make "K" the subject sqrt{} C^2-ak=b

OpenStudy (anonymous):

\[\sqrt{} C^2-ak=b\]

OpenStudy (sasogeek):

\[\sqrt{C ^{2}-ak}=b\] or \[\sqrt{C ^{2}}-ak=b\] or\[\sqrt{C}^{2}-ak=b\] which one?

OpenStudy (anonymous):

the one you typed first..

OpenStudy (sasogeek):

ok so what you want to do is get rid of the square root first... do you know how to do that?

OpenStudy (anonymous):

yeah by squaring the numbers in the right. and then..can u please do the working plus the answer

OpenStudy (anonymous):

not numbers letters*

OpenStudy (anonymous):

thanks alot bro...u rock...

OpenStudy (sasogeek):

;)

OpenStudy (sasogeek):

anytime :)

OpenStudy (anonymous):

'subject'?